OpenStudy (anonymous):

Step-by-Step How do i solve x²-36=0

jhonyy9 (jhonyy9):

us a2 -b2 =(a-b)(a+b)

jhonyy9 (jhonyy9):

x2 -36=(x-6)(x+6)

jhonyy9 (jhonyy9):

so x-6=0 x=6 and x+6=0 x=-6

OpenStudy (anonymous):

Another intuitive way Say that: \[x^{2}\] = 36 Take the square root of each side and you get that x = 6, -6

jhonyy9 (jhonyy9):

yes is right too